Jim is selling basketball tickets for a game. Adult tickets are $5 and $3 for student tickets. After 2 hours he sells 50 tickets and earns $180. How many adult tickets how many student tickets did he sell?
Answer by checkley77(12844)

5x+3(50-x)=180
5x+150-3x=180
2x=180-150
2x=30
x=30/2
x=15 adult tickets sold.
50-15=35 student tickets sold.
Proof:
5*15+35*3=180
75+105=180
180=180